# Automated Statistical Model Discovery with Language Models

### v1

@@ Results
In Figure 5 (left), we plot the predictions obtained from integrating the learned dynamics for an ODE proposed by the LM, the training data points generated from the true dynamics, and the predictions from standard Lotka-Volterra. For all ODEs, we fit free parameters to the training data via gradient descent. The grey region indicates the extrapolation region.
The LM can significantly improve upon the standard Lotka-Volterra model by introducing corrections to the dynamics (Figure 5). The standard Lotka-Volterra model cannot capture the decreasing amplitude in the data; furthermore, there is a slight phase shift relative to the training datapoints. In contrast, the LM identifies an ODE that captures these properties and extrapolates accurately. In Figure 5 (right), we compare these programs against a neural ODE baseline, and a hybrid neural ODE baseline that incorporates a multiplicative correction (parameterized by an MLP) to the predator-prey interaction term in the predator equation. See Section C of the Appendix for details on these baselines. For LM variations, we report the average test MAE across three runs. All LM variations outperform the baselines.
In Figure 6, we present code snippets corresponding to representative programs proposed in the constrained and unconstrained variations. These snippets show how natural language constraints can guide the LM towards more flexible models that retain interpretability. For the variation with no natural language constraints, the LM takes a purely empirical approach. In particular, the LM adjusts the scaling terms on the additive correction term parameterized by an MLP (Figure 8). For the constrained variation, the LM proposes a hybrid approach integrating the neural approach in the prompt with classic models in the literature; importantly, even though the LM is asked to balance interpretability and flexibility, the LM still identifies programs that outperform the Neural ODE and standard Lotka-Volterra baselines. One approach the LM proposes is an extension of the Rosenzweig and MacArthur model with a Holling’s type II functional response (Rosenzweig & MacArthur, 1963) to allow a static parameter to depend dynamically on the prey density: the LM models the handling time, or the time a predator spends “processing” a prey, as a nonlinear function of the prey density via an MLP. These results show that natural language constraints drive LMs towards models that balance flexibility and interpretability.
@@ 4 Conclusion
We introduced a method for leveraging LMs for automated model discovery. We showed our method can identify models that perform favorably against strong baselines. We also studied how domain knowledge and natural language constraints influence the programs proposed by LMs. Altogether, our results highlight the compelling advantages of LM driven statistical model discovery. Although we focused on the static data regime in this work, a promising future direction is to explore settings where the LM designs experiments to collect more data.
5 Acknowledgements
@@ Appendix B Stan Experiments
Language model produced pymc code snippet for modeling Peregrine population count
1def dugongs_gen_model():
2  # Priors for von Bertalanffy growth function parameters
3  L_inf = pm.Uniform(’L_inf’, lower=0, upper=3) # Asymptotic length, bounded by observed data
4  k = pm.Uniform(’k’, lower=0, upper=1) # Growth coefficient, bounded by reasonable values
5  t0 = pm.Uniform(’t0’, lower=-5, upper=5) # Theoretical age at zero length
6  # Expected length at each age using von Bertalanffy growth function
7  expected_length = L_inf * (1 - pm.math.exp(-k * (ages - t0)))
8  # Likelihood for observed lengths
9  sigma = pm.Uniform(’sigma’, lower=0, upper=1) # Standard deviation of observed lengths around the mean
10  y_obs = pm.Normal(’y_obs’, mu=expected_length, sigma=sigma, observed=lengths)
@@ Appendix C Lotka-Volterra
1def no_warm_start(t, y, _coeffs): 2  # Hypothesis 1: Logistic growth for prey with carrying capacity kappa 3  # Hypothesis 2: Saturation effect in predation rate with parameter psi 4  db_dt = alpha * b * (1 - b / kappa) - beta * b * c / (1 + psi * b) 5  dc_dt = -gamma * c + delta * b * c - epsilon * c**2 1def warm_start_constrained(t, y, _mlp): 2  inputs = jnp.array([b])  # Input is prey density (bunny population) 3  # Neural network learns a modifier for the handling time based on prey density 4  handling_time_modifier = _mlp(inputs) 5  # Modulate the predation rate beta in a non-linear manner according to handling time 6  db_dt = alpha * b - beta * b * c / (1 + handling_time_modifier[0] * b) 7  dc_dt = -gamma * c + delta * b * c / (1 + handling_time_modifier[0] * b) 1def warm_start_unconstrained(t, y, _mlp): 2  inputs = jfu.ravel_pytree((b, c))[0] 3  mlp_output = _mlp(inputs) 4  # Fine-tuned scaling of the MLP output to match the amplitude of data more closely 5  db_dt = alpha * b - beta * b * c + 0.02 * mlp_output[0] # Reduced scaling for bunnies 6  dc_dt = -gamma * c + delta * b * c + 0.06 * mlp_output[1] # Increased scaling for cats
Figure 8: LM can propose corrections to ODEs (top) In the no warm-start variation, the LM introduces corrections informed by knowledge of predator-prey models (carrying capacity, predation saturation) (middle) When prompted to introduce neural networks in an interpretable way, LM proposes to make the handling time parameter depend non-linearly on the prey density, extending the Holling’s type II response, a traditional approach to modeling predation saturation. (bottom) When prompted to introduce neural networks without constraints, LM introduces additive MLP parameterized corrections and adjusts the scaling factors.

### v2

@@ Results
In Figure 4 (left), we plot the predictions obtained from integrating the learned dynamics for an ODE proposed by BoxLM, the training data points generated from the true dynamics, and the predictions from the standard Lotka-Volterra model. We fit free parameters to the training data via gradient descent. The grey region indicates the extrapolation region.
BoxLM can significantly improve upon the standard Lotka-Volterra model by introducing corrections to the dynamics (Figure 4). The standard Lotka-Volterra model cannot capture the decreasing amplitude in the data; furthermore, there is a slight phase shift relative to the training datapoints. In contrast, BoxLM identifies an ODE that captures these properties and extrapolates accurately. In Figure 4 (right), we compare these programs against a neural ODE baseline, and a hybrid neural ODE baseline that incorporates a multiplicative correction (parameterized by an MLP) to the predator-prey interaction term in the predator equation. See Section C of the Appendix for details on these baselines. For the LM variations, we report the average test MAE across three runs. In Figure 4, we see that all BoxLM variations outperform the baselines.
In Figure 5, we present code snippets corresponding to representative programs proposed in the constrained and unconstrained variations. These snippets show how natural language constraints can guide BoxLM towards more flexible models that retain interpretability. For the variation with no natural language constraints (WS-No Constraint), BoxLM takes a purely empirical approach. In particular, BoxLM adjusts the scaling terms on the additive MLP correction term. For the WS-Constraint variation, BoxLM proposes a hybrid approach integrating the neural approach in the prompt with classic models in the literature; importantly, even though BoxLM is asked to balance interpretability and flexibility, BoxLM still identifies programs that outperform the neural ODE and standard Lotka-Volterra baselines. One approach BoxLM proposes is an extension of the Rosenzweig and MacArthur model with a Holling’s type II functional response (Rosenzweig & MacArthur, 1963) to allow a static parameter to depend dynamically on the prey density: BoxLM models the handling time, or the time a predator spends “processing” a prey, as a nonlinear function of the prey density via an MLP. These results show how we can use natural language to drive BoxLM towards models that balance flexibility and interpretability.
@@ 4 Conclusion
We introduced a method for leveraging LMs for automated model discovery. Our method can identify models that perform favorably against strong baselines and improve upon expert models. We also studied how domain knowledge and natural language constraints influence our system. Altogether, our results highlight the compelling advantages of LM-driven statistical model discovery.
Our work has important limitations that motivate future research. First, we focused on modeling static datasets. An interesting direction could be leveraging LMs for active data collection. Second, since our tasks were restricted to one-dimensional datasets, simple model criticism statistics were sufficient and therefore decided in advance (residuals, posterior predictive mean). Another interesting future direction could be fully automating the criticism step. Finally, while in-context learning was effective in our tasks, we could explore finetuning techniques for training a language model to produce better probabilistic programs.
Figure 6: Round-to-round improvement. Model revision leads to improvements on average. (Improvement is not necessarily monotonic for a given dataset and run.) (top) ELPD LOO score vs. round for Stan experiments. We normalize the ELPD LOO scores across programs proposed for 3 rounds of Box’s loop and average across datasets for the no metadata condition. Larger is better and error bars correspond to standard error. (bottom) Squared error vs round for LV experiment. We report the squared error averaged across three different runs, for the warm-start, no constraint condition; smaller is better.
Acknowledgements
@@
We thank Eric Zelikman and Neil Band for detailed feedback on this paper. We also thank Omar Shaikh and Jensen Gao for helpful discussions.
Impact Statement
This paper presents work whose goal is to partially automate statistical modeling. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix B Stan Experiments
Language model produced pymc code snippet for modeling Peregrine population count 1def dugongs_gen_model(): 2 # Priors for von Bertalanffy growth function parameters 3 L_inf = pm.Uniform(’L_inf’, lower=0, upper=3) # Asymptotic length, bounded by observed data 4 k = pm.Uniform(’k’, lower=0, upper=1) # Growth coefficient, bounded by reasonable values 5 t0 = pm.Uniform(’t0’, lower=-5, upper=5) # Theoretical age at zero length 6 # Expected length at each age using von Bertalanffy growth function 7 expected_length = L_inf * (1 - pm.math.exp(-k * (ages - t0))) 8 # Likelihood for observed lengths 9 sigma = pm.Uniform(’sigma’, lower=0, upper=1) # Standard deviation of observed lengths around the mean 10 y_obs = pm.Normal(’y_obs’, mu=expected_length, sigma=sigma, observed=lengths)
@@ Appendix C Lotka-Volterra
Appendix D Failure rates of GPT-4 V proposed programs
Model Percent successfully scored
GPT-4 textual dataframe 78%
GPT-4 vision only 70%
GPT-3.5 76%
Table 2: Percentage of pymc programs that we can successfully perform inference in for Stan experiments.
Appendix E Additional ODE results
Figure 8: Modeling nonlinear ODEs. LM can introduce polynomial corrections to simple harmonic oscillator (SHO) that provide a better fit. Grey region indicates extrapolation region.
We evaluated on three additional ODEs (nonlinear oscillators) from this github repository pysindy. These ODEs include a Duffing ODE, Van Der Pol oscillator, and a Cubic Harmonic Damped oscillator. We give BoxLM a simple harmonic oscillator to start and ask it to introduce polynomial corrections to improve the fit. Consistent with Section 3.3, our approach can generally improve upon a baseline ODE and extrapolate accurately into the test region.
Appendix F Visual Interface/GPT-3.5 Ablations
Context length limits are a potential limitation if we provide the dataset in textual format in the prompt. We therefore experiment with a visual-only variation. We remove all textual representations of datasets and model criticism statistics from the prompt and only provide visual plots of these datasets and statistics. We show this visual-only variation does not harm the performance relative to textual variation and should not suffer as much from context length limits as the dataset grows larger. We also show we can obtain similar results with GPT-3.5 Turbo, which is significantly less costly than GPT-4.
Dataset GPT-4 Text GPT-4 Visual Only GPT-3.5 Expert
Eight schools -30.17 -30.40 -30.44 -30.70
Dugongs 22.61 23.76 21.01 22.52
Surgical -37.36 -38.54 -42.2 -40.29
Peregrine -164.69 -143.45 -161.14 -142.19
Table 3: Vision interface and model type ablations. Comparison of GPT-4 with textual representation of data and model criticism statistics in prompt, GPT-4 with only visual representations (GPT-4 Visual Only) of data and model criticism statistics (e.g., only plots), and GPT 3.5 Turbo against expert programs. The visual-only variation (GPT-4 Visual-Only) performs comparably to the textual variation (GPT-4 Text )and outperforms the textual variation on the Peregrine dataset. GPT-3.5 performs slightly worse on some datasets but comparably on most. We report the expected predictive log density (LOO).
Appendix G State-space model hidden state update experiments
We present the same results from the main text but take a state-space update inspired approach to computing the natural language criticism $h^{t}$.
Dataset BoxLM+ State Space BoxLM State Space Periodic AS SM N-BEATS
Air 0.04 0.34 0.15 0.19 0.06 0.22
Beer 0.07 0.41 0.15 0.06 0.05 0.02
Heart 0.20 0.25 0.20 0.21 0.21 0.07
Milk 0.06 0.11 0.10 0.11 0.09 0.04
Wine 0.17 0.17 0.21 0.13 0.18 0.17
Wool 0.24 0.15 0.19 0.23 0.13 0.18
Figure 9: Test set performance of BoxLM with state-space updated on time series datasets. Performance of BoxLM system with state space update for model criticism. Comparison of BoxLM test mean absolute error (MAE) against Automatic Statistician using greedy search (AS), spectral mixture kernel (SM), periodic kernel (Periodic), and N-BEATS. BoxLM+ searches over an augmented kernel space. We bold the best and underline the second best among the GP methods.
Dataset Expert LM
Eight schools -30.70 -30.17
Eight schools sim -18.09 -18.39
Eight schools sim no metadata -18.09 -18.90
Dugongs 22.52 22.61
Dugongs sim 50.04 57.4
Dugongs sim no metadata 50.04 26.68
Surgical -40.29 -37.36
Surgical sim -39.80 -38.45
Surgical sim no metadata -39.80 -58.72
Peregrine -142.19 -164.69
Peregrine sim -130.48 -177.15
Peregrine sim no meta -130.48 -127.32
Table 4: Comparison of BoxLM with state-space update programs against expert programs We perform this comparison across four different datasets and two different ablations that replace observations with synthetic observations and remove all metadata. We report the expected predictive log density estimated via leave-one-out cross validation. We bold statistically significant differences and underline non-significant differences.
Figure 10: Correcting misspecified Lotka-Volterra dynamics (BoxLM with state-space updates). BoxLM can introduce corrections to standard Lotka-Volterra dynamics (no warm-start) and a hybrid neural ODE approach (warm-start) that outperform several baselines. Test MAE of LM models (No-WS, WS-Constraint, and WS-No Constraint) compared to the standard Lotka-Volterra model LV, a Neural ODE, and a hybrid Neural ODE model with a multiplicative correction to the prey-predator dynamics (Hybrid).